OpenStudy (anonymous):

How do I find the length of side a in the right triangle, if b=6 and c= 10? I need to find the exact answer, using radicals as needed.

jimthompson5910 (jim_thompson5910):

\[a^2+b^2=c^2\] \[a^2+6^2=10^2\] \[a^2+36=100\] \[a^2=100-36\] \[a^2=64\] \[a=\sqrt{64}\] \[a=8\]